OpenStudy (anonymous):

What is the sum of the solutions of ? 2|x-1|-4=-2 A. 0 B. 1 1/2 C. 2 D. -2

OpenStudy (anonymous):

2(x−1)−4=−2 I will try to solve your equation. 2x−6=−2(Simplify both sides of the equation) 2x−6+6=−2+6(Add 6 to both sides) 2x=4 (2x) 2 = 4 2 (Divide both sides by 2) x=2

OpenStudy (anonymous):

Is that one part completed? Like are we supposed to solve the opposite of that too o. o

OpenStudy (anonymous):

No, the answer is C

OpenStudy (anonymous):

Is that - 4 or negative4

OpenStudy (anonymous):

Because I solved using - 4

OpenStudy (anonymous):

Oooh... I suppose. Thank you and it's - 4

OpenStudy (raden):

2|x-1|-4=-2 2|x-1| = -2 + 4 2|x-1| = 2 |x-1| = 2/2 |x-1| = 1 x - 1 = 1 x_1 = 2 or x - 1 = -1 x_2 = 0 the sum of x's = 2 + 0 = 2

OpenStudy (anonymous):

Ok, goody! :D

OpenStudy (anonymous):

I already solved it

OpenStudy (raden):

but you have one x, while we should we have 2 x's. but the last answer is correct :)

OpenStudy (anonymous):

Thank you both c:

OpenStudy (raden):

welcome